OpenStudy (anonymous):

A cheetah ran 300 feet in 2.92 seconds. What was the cheetah’s average speed in miles per hour?

OpenStudy (anonymous):

300 divided by 2.92

OpenStudy (john_es):

And then apply unit conversion.

OpenStudy (calculusxy):

Since you are comparing with the feet to the seconds, set this up into a division equation because you have to find the mph. So just divide 300/2.92=102.7339...~102.73.

OpenStudy (john_es):

If you use the equivalence, \[1 \text{feet}=0.000189 \text{miles}\]\[2.29s=2.92/3600\ \ \text{hour}\] Then, \[300\cdot0.000189/(2.92/3600)=69.9\ \ \text{miles/hour}\]